# Decision problems on geometric tilings

### v1

@@ Abstract
We study decision problems on geometric tilings. First, we study a variant of the Domino problem where square tiles are replaced by geometric tiles of arbitrary shape. We show that, under some weak assumptions, this variant is undecidable regardless of the shapes, extending the results of [11] on rhombus tiles. This result holds even when the geometric tiling is forced to belong to a fixed set. Second, we consider the problem of deciding whether a geometric subshift has finite local complexity, which is a common assumption when studying geometric tilings. We show that it is undecidable even in a simple setting (square shapes with small modifications).
@@ 1 Introduction
The undecidability of the domino problem [5] is a founding result of the theory of symbolic tilings: given a set of forbidden patterns, is it possible to tile the plane while avoiding this set of forbidden pattern? The domino problem has subsequently been studied on many structures: many groups [2], the hyperbolic plane [17, 8], on tilings by rhombuses by the two first authors [11], inside other symbolic tilings [3]. In this article we focus on a fusion of the two models, symbolic geometric tilings: they are geometric tilings with added decorations and a finite number of forbidden patterns on these decorations (precise definitions are in subsection 2.4).
Our first problem of interest is the domino problem, as studied in [11] for rhombus tilings: fixing a nonempty set of geometric tilings $X$, given as inputs a set decorations on the shapes of $X$ and a set of forbidden patterns, can we tile the plane with shapes of $X$ decorated in such a way that they avoid all forbidden patterns? We show that this problem is always undecidable under weak assumptions on the shapes and $X$.
@@
The second problem we study is the following: given a geometric tileset, is it decidable whether it has FLC?
@@ 2.1 Computability
A decision problem $A:\mathbb{N}\subseteq\{0,1\}$ is said to be computable when there exists a Turing machine $M$ that halts on all inputs $x\in\mathbb{N}$, and which outputs $1$ iff $A(x)=1$. The canonical undecidable problem is the halting problem, denoted $\mathbf{HP}$: given (the number corresponding to) a Turing machine as an input, does it halt on an empty input tape?
The halting problem is recursively enumerable: there is a Turing machine which never halts and outputs every input that map to $1$ (in this case, every Turing machine that halts). Moreover, it is complete for the class of recursively enumerable problems under many-one reduction, defined as follows: we say that $A\leq_{m}B$ if there exists a total computable function $f$ such that $B(x)=1\Leftrightarrow A(x)=1$. The completeness of the halting problem means that for any recursively enumerable problem $A$, $A\leq_{m}\mathbf{HP}$.
@@ 2.2 Symbolic tilings
$d(x,y)=2^{-min\{\left\lVert z\right\rVert\mid z\in\mathbb{Z}^{2},\leavevmode\nobreak\ x(z)\neq y(z)\}}\text{.}$
@@ 2.3 Geometric tilings
We choose a distance between geometric tilings that generalises the usual distance on $\mathbb{Z}^{2}$ configurations. Two tilings $T_{1}$ and $T_{2}$ are $\varepsilon$-close if they are identical on a disk of radius $\varepsilon^{-1}$ up to a translation of norm at most $\varepsilon$. See [18, 16] for full details.
@@
We introduce restriction on the geometric tilings on which our main result applies:
Definition 2.8 (Potatoes).
A shape is a potato if its boundary contains at most countably many maximal horizontal and vertical segments (that are not included in a larger segment).
Our results hold if this assumption applies for any two fixed non-colinear directions, but this makes the proofs more tedious.
Remark 2.9 (Smooth shapes, polygons).
As argued in [14], any geometric tiling by topological disks is combinatorially equivalent to a polygonal tiling (where all shapes are polygons), though some properties of the tilings such as self-similarity might be lost in this equivalence. We use the very weak potato assumption for technical reasons.
@@ 2.4 Symbolic-geometric tiling and the domino problem
The case where $\mathcal{T}$ contains only a unit square and $X$ is the subshift of edge-to-edge square tilings (containing a single point up to translation) corresponds to the classical domino problem. This definition, as well as an equivalent of the following result, appeared in [11] in the particular case of rhombus shapes.
@@ 3 The domino problem is undecidable on every set of tilings
Let $X$ be a non-empty set of tilings on potato shapes with finite local complexity. The domino problem on $X$ is $\Pi_{1}^{0}$-hard and therefore undecidable.
@@ 3.1 Coloured tiles
We assume (by rescaling if necessary) that each shape in $\mathcal{T}$ strictly contains a unit square and do not have in its boundary two vertical or horizontal segments at integer distance from each other, using the potato assumption. For each shape, we pick an arbitrary origin point from which we draw a square grid of width $\tfrac{1}{4}$, as illustrated in Fig. 5 (we pick a vertex as origin point in the figures). We call checkered tiles these shapes with the superimposed $\tfrac{1}{4}$-grid.
Figure 5: The $\tfrac{1}{4}$ square grid superimposed on the square and triangle shape of $X_{\triangle}$ forming the checkered square-triangle tiles.
@@
We note $\mathbf{T}_{0}$ the set of placeholder tiles.
@@
For each placeholder tile $\mathbf{t}_{0}$ in $\mathbf{T}_{0}$, we define a coloured tile $\mathbf{t}$ for each $\Sigma$-labelling of the coding cells. We note $\mathbf{T}$ the set of coloured tiles obtained from this process. As the shapeset is finite, the size of the shapes is bounded; assuming each tile in $\mathbf{T}_{0}$ contains at most $n$ coding cells, it defines at most $|\Sigma|^{n}$ coloured tiles, so $\mathbf{T}$ is finite.
@@
As there are 5 coding cells and 5 colours, the total number of coloured tiles for this placeholder is $5^{5}$, most of which are later forbidden by $G_{F}$. The five rightmost tiles are allowed and the three leftmost tiles are later forbidden as they contain a forbidden pattern or force a forbidden pattern.
@@ 3.2.1 Support of the forbidden patterns
Let $D_{max}$ be the maximal diameter of the shapes in $\mathcal{T}$. Given a $r$-pattern $c$, we may assume that its center is 00 up to translating the pattern. Since a tile in $c$ either intersects $B(0,r)$ or fills a hole left by tiles that intersect $B(0,r)$, it follows that $c$ is contained in $B(0,r+D_{max})$. By finite local complexity, the set of labelled $r$-patterns is finite.
@@ 3.2.2 Grid structure
Grid-like labelled patterns exist on any geometric pattern. Indeed, take a pattern $P_{0}$ of unlabelled shapes and the induced pattern $P_{1}$ of checkered tiles. Given a cell $c$ of a checkered tile, the set of positions $(x_{0},y_{0})$ in the interior of $c$ such that the grid $g_{x_{0},y_{0}}$ contains a tile triple point or a segment intersection with the tile boundaries or the placeholder grid boundary is countable (since shapes are potatoes, they have at most countably many maximal horizontal and vertical segments). By a simple counting argument, there are uncountably many positions $(x_{0},y_{0})$ in the interior of $c$ whose grid satisfies conditions 1 and 2, and for each such position it is straightforward to check that there is a unique choice of placeholder tile for each checkered tile that satisfies condition 3: since each tile contains a vertex of the grid, this forces the position of a coding cell.
@@ 3.3 Equivalence with the original domino problem
which implies that $f(1,1)=v$ (they are vertices on a unit grid). Since $p^{\prime}$ is in a coding cell and is at distance $<\tfrac{1}{4}$ from a coding cell with colour $c_{1,1}$, the two cells have the same colour by $G_{coding}$. It follows that the grid centered on $f(0,1)$sees" the pattern $c_{(0,1),(1,1)}$, so the $D_{max}$-pattern around $f(0,1)$ belongs to $G_{F}$.
@@ 3.4 Hidden quasi-isometries
In the case of finitely presented groups it is known that the undecidability of the domino problem is a quasi-isometry invariant [6]. In particular, the domino problem on any Cayley graph that is quasi-isometric to $\mathbb{Z}^{2}$ is undecidable. However the proof uses heavily the group structure and we do not claim that the undecidability of the domino problem is a quasi-isometry invariant on general (infinite) graphs. There are other quasi-isometry invariant group properties related with the existence of strongly aperiodic subshifts [4].
Adjacency graphs are very far from being Cayley graphs: say two vertices are equivalent if they have the same neighbourhoods for all radii. Cayley graphs are transitive (all vertices are equivalent) while adjacency graphs are not even quasi-transitive (there may be infinitely many equivalence classes). Still, our proofs seem to hint that quasi-isometries play a role even in this context.
@@ 4 Finite local complexity is undecidable for geometric tilings
a shapeset $\mathcal{T}$.
@@
The FLC problem for purely geometric subshifts is $\Pi_{1}^{0}$-hard and therefore undecidable.
Proof 4.3.
We proceed by reduction to the halting problem for Turing machines. For each Turing machine $M$, we build a shapeset $\mathcal{T}$ such that the purely geometric subshift is not empty and has finite local complexity iff $M$ does not halt on the empty tape.
We start from $1\times 1$ square tiles decorated with the Robinson [19] decorations as modified by Gurevich and Koryakov [10]: this tileset tiles the plane periodically iff the Turing machine encoded in it halts and aperiodically otherwise, as shown in Figure 12. Robinson’s tileset can be defined either by purely geometric or symbolic tilings: see Fig. 13. Gurevich and Koryakov’s tileset can also be made purely geometric: see Figs. 14 and 15.
@@
Figure 13: A purely geometric presentation of Robinson tiling. Decorations are shown for clarity only, and the computing layer is not represented.
@@
(a) Schematics of the embedding of computation in a Robinson or Gurevich-Koryakov tile.
(b) Two examples of computing tiles.
Left: from state $q_{1}$, reading a 00, the machine writes a 00, goes in state $q_{2}$ and does not move.
Right: from state $q_{2}$, reading a 00, the machine writes a $1$, stays in state $q_{2}$ and moves to the right.
@@
If $M$ does not halts, any individual tiling in the subshift $X$ has finite local complexity (even though $X$ doesn’t), since it contains at most one shear line with a given offset which yields only finitely many $2$-shapes patterns. The construction can be tweaked so that individual tilings do not have FLC in this case by duplicating our shapeset into squares of size 1 and squares of size $\varphi=(1+\sqrt{5})/2$ (or any irrational number). Tiles inside a finite or infinite green square synchronise so that only one size is used. However, when there is a shear line (in the case where the machine does not halt), the two half-planes may use different shapesets, as illustrated on Fig. 19. The offset between adjacent tiles across the shear lines are the multiples of $\varphi$ modulo $1$ (plus an initial offset) which are dense in the interval $[0,1]$. So such a tiling does not have finite local complexity.
@@ Appendix A Aperiodic Wang tilesets from aperiodic geometric tilesets
This construction was first used in [9, §11.1] for Kite-and-Darts Penrose tilings to produce an aperiodic Wang tileset of $32$ tiles, and left as exercise (11.1.2) for rhombus Penrose tilings to produce an aperiodic Wang tileset of $24$ tiles; full details on this latter case can be found in [12, §5]. We give an overview of this construction, in particular to shed light on geometric and combinatorial conditions that make it work.
We present the construction for rhombus tilings though it can also work for e.g. the Kites-and-Darts. Let $\mathcal{T}$ be a finite set of rhombus shapes and $\mathbf{T}$ a finite labelled tileset on $\mathcal{T}$. Let $X$ be an edge-to-edge $\mathbf{T}$ nearest-neighbour subshift of finite type, that is, defined by finitely many forbidden patterns made of two edge-adjacent tiles.
@@
  • the size of $t_{0}$-chain-squares is bounded, so there are finitely many different $t_{0}$-chain-squares in $X$. A formal proof of a very similar statement is given in e.g. [11, Lemma 21].
Figure 22: A fragment of labelled Penrose rhombus tiling with occurences of a tile highlighted.

### v2

@@ Abstract
We study decision problems on geometric tilings. First, we study a variant of the Domino problem where square tiles are replaced by geometric tiles of arbitrary shape. We show that this variant is undecidable regardless of the shapes, extending the results of [15] on rhombus tiles. This result holds even when the geometric tiling is forced to belong to a fixed set. Second, we consider the problem of deciding whether a geometric subshift has finite local complexity, which is a common assumption when studying geometric tilings. We show that this problem is undecidable even in a simple setting (square shapes with small modifications).
@@ 1 Introduction
The undecidability of the domino problem [7] is a founding result of the theory of symbolic tilings: given a set of forbidden patterns, is it possible to tile the plane while avoiding this set of forbidden pattern? The domino problem has subsequently been studied on many structures: many groups [2], the hyperbolic plane [24, 19, 11], tilings by rhombuses by the first two authors [15], other symbolic tilings [3], self-similar fractal structures [5], etc. In this article we focus on a fusion of the two models, symbolic-geometric tilings: they are geometric tilings with colored local rules, that consist in adding decorations and a finite number of forbidden patterns on these decorations (precise definitions can be found in subsection 2.4). Colored local rules have been used to capture properties of geometric tilings, such as when they are generated by a substitution [12] or planar [9].
Our first problem of interest is the domino problem, as studied in [15] for rhombus tilings: fixing a nonempty set of geometric tilings $X$, given as input some colored local rules on the shapes of $X$, can we tile the plane with shapes of $X$ decorated in such a way that they avoid all forbidden patterns? We show that this problem is always undecidable when the shapes are compact and simply connected. The proof can be seen as representing a quasi-isometry between $\mathbb{Z}^{2}$ and the adjacency graph of a geometrical tiling, in order to reduce the domino problem on $\mathbb{Z}^{2}$ which is known to be undecidable. We discuss the role of quasi-isometries and related results in Section 3.4.
@@
The second problem we study is the following: given a geometric tileset, is it decidable whether it has FLC? We show that having FLC is a $\Sigma^{0}_{1}$-complete problem and is thus undecidable.
@@ 2.1 Computability
A decision problem $A:\mathbb{N}\subseteq\{0,1\}$ is said to be computable when there exists a Turing machine $M$ that halts on all inputs $x\in\mathbb{N}$, and which outputs $1$ iff $A(x)=1$. The canonical undecidable problem is the halting problem, denoted $\mathbf{HP}$: given (the number corresponding [^2] Using the usual enumeration of Turing machines. to) a Turing machine as an input, does it halt on an empty input tape?
The halting problem is recursively enumerable: there is a Turing machine which never halts and outputs every input that maps to $1$ (in this case, every Turing machine that halts). Moreover, it is complete for the class of recursively enumerable problems under many-one reduction, defined as follows: we say that $A\leq_{m}B$ if there exists a total computable function $f$ such that $B(f(x))=1\Leftrightarrow A(x)=1$. The completeness of the halting problem means that for any recursively enumerable problem $A$, $A\leq_{m}\mathbf{HP}$.
@@ 2.2 Symbolic tilings
$d(x,y)=2^{-min\{\left\lVert z\right\rVert\mid z\in\mathbb{Z}^{2},\penalty 10000\ x(z)\neq y(z)\}}\text{.}$
@@ 2.3 Geometric tilings
Remark 2.5 (Shapes and polygons).
As argued in [20], any geometric tiling by topological disks is combinatorially equivalent to a polygonal tiling (where all shapes are polygons), though some properties of the tilings such as self-similarity might be lost in this equivalence. We write our results for topological disks as it does not make the proofs significantly more complicated; the next proposition is the only technical property that we use.
Proposition 2.6.
The boundary of a shape $t$ contains at most countably many maximal segments (that are not included in a larger segment).
Proof 2.7.
A shape $t$ is, by assumption, homeomorphic to a unit disk. The images under this homeomorphism of maximal segments in the boundary of $t$ are disjoint connected subsets of the unit circle, that is, intervals, that each contain some point of the form $e^{ri\pi}$ for $r\in\mathbb{Q}$; therefore there cannot be more than countably many of them.
@@
We choose a distance between geometric tilings that generalises the usual distance on $\mathbb{Z}^{2}$ configurations. Two tilings $T_{1}$ and $T_{2}$ are $\varepsilon$-close if they are identical on a disk of radius $\varepsilon^{-1}$ up to a translation of norm at most $\varepsilon$. See [25, 27, 23] for full details.
@@ 2.4 Symbolic-geometric tiling and the domino problem
Forbidden patterns on decorated shapes have also been called colored local rules in the litterature. The case where $\mathcal{T}$ contains only a unit square and $X$ is the subshift of edge-to-edge square tilings (containing a single point up to translation) corresponds to the classical domino problem. This definition, as well as an equivalent of the following result, appeared in [15] in the particular case of rhombus shapes.
@@ 3 The domino problem is undecidable on every set of tilings
Let $X$ be a non-empty set of tilings with finite local complexity. The domino problem on $X$ is $\Pi_{1}^{0}$-hard and therefore undecidable.
@@ 3.1 Coloured tiles
We assume (by rescaling if necessary) that each shape in $\mathcal{T}$ strictly contains a unit square and does not have in its boundary two vertical or horizontal segments at integer distance from each other, using Proposition 2.6. For each shape, we pick an arbitrary origin point from which we draw a square grid of width $\tfrac{1}{4}$, as illustrated in Fig. 5 (we pick a vertex as origin point in the figures). We call checkered tiles these shapes with the superimposed $\tfrac{1}{4}$-grid.
The choice to superimpose a grid of width $\tfrac{1}{4}$, together with the choice of coding cells below, ensures that: any labelled tile contains a coding cell (Fig. 6), adjacent tiles have coherent colourings of their common coding cells (Fig. 10), and in any tiling the coding cells are arranged on a valid $\mathbb{Z}^{2}$ grid (Section 3.2.2). This grid could be replaced with any finer grid (for example of width $\tfrac{1}{k}$ with $k\geq 4$): in that case we might change the choice of coding cells.
Figure 5: The $\tfrac{1}{4}$ square grid superimposed on the square and triangle shape of $X_{\triangle}$ forming the checkered square-triangle tiles. Recal that we defined the square and triangle shapes with sides $\tfrac{5}{2}$.
@@
We define $\mathbf{T}_{0}$ the set of placeholder tiles.
@@
We now decorate each placeholder tile $\mathbf{t}_{0}$ in $\mathbf{T}_{0}$, using the alphabet of the SFT $(A,F)$: we define a coloured tile $\mathbf{t}$ for each possible $A$-labelling of the coding cells. We denote by $\mathbf{T}$ the set of coloured tiles obtained from this process. As the shapeset is finite, the diameter of the shapes is bounded and therefore they contain a bounded number of coding cells; assuming each tile in $\mathbf{T}_{0}$ contains at most $n$ coding cells, it defines at most $|A|^{n}$ coloured tiles, so $\mathbf{T}$ is finite.
@@
As there are 5 coding cells and 5 colours, the total number of coloured tiles for this placeholder is $5^{5}$, most of which are later forbidden by $G_{F}$. The five leftmost tiles are allowed and the three rightmost tiles are later forbidden as they contain a forbidden pattern or force a forbidden pattern.
@@ 3.2.1 Support of the forbidden patterns
Let $D_{max}$ be the maximal diameter of the shapes in $\mathcal{T}$. Given a $r$-pattern $c$, we may assume that its center is 0 up to translating the pattern. Since a tile in $c$ either intersects $B(0,r)$ or fills a hole left by tiles that intersect $B(0,r)$, it follows that $c$ is contained in $B(0,r+D_{max})$. By finite local complexity, the set of labelled $r$-patterns is finite.
@@ 3.2.2 Grid structure
Grid-like labelled patterns exist on any geometric pattern. Indeed, take a pattern $P_{0}$ of unlabelled shapes and the induced pattern $P_{1}$ of checkered tiles. Given a cell $c$ of a checkered tile, the set of positions $(x_{0},y_{0})$ in the interior of $c$ such that the grid $g_{x_{0},y_{0}}$ contains a tile triple point or a segment intersection with the tile boundaries or the placeholder grid boundary is countable (by Proposition 2.6, shapes have at most countably many maximal horizontal and vertical segments). By a simple counting argument, there are uncountably many positions $(x_{0},y_{0})$ in the interior of $c$ whose grid satisfies conditions 1 and 2, and for each such position it is straightforward to check that there is a unique choice of placeholder tile for each checkered tile that satisfies condition 3: since each tile contains a vertex of the grid, this forces the position of a coding cell.
@@ 3.3 Equivalence with the original domino problem
which implies that $f(1,1)=v$ (they are vertices on a unit grid). Since $p^{\prime}$ is in a coding cell and is at distance $\tfrac{1}{4}$ from a coding cell with colour $c_{1,1}$, the two cells have the same colour by $G_{coding}$. It follows that the grid centered on $f(0,1)$ "sees" the pattern $c_{(0,1),(1,1)}$, so the $D_{max}$-pattern around $f(0,1)$ belongs to $G_{F}$.
@@ 3.4 Hidden quasi-isometries
In the case of finitely presented groups it is known that the undecidability of the domino problem is a quasi-isometry invariant [8], in addition to other properties such as the existence of strongly aperiodic subshifts [6]. In particular, the domino problem on any Cayley graph that is quasi-isometric to $\mathbb{Z}^{2}$ is undecidable. However, adjacency graphs for geometric tilings are very far from Cayley graphs; for example, they are not (vertex-)transitive.
Shortly after the present paper was submitted, Barbieri and Bitar [4] extended these invariance results under quasi-isometry beyond Cayley graphs to structures called blueprints; see in particular Section 6 in op.cit that relate to geometric tilings. While their results are more general than Theorem 3.1 on most aspects (they apply to all groups and not only $\mathbb{Z}^{2}$, for example), they only cover the case where the set of geometric tilings is a subshift as of the currently available version [^4] Following personal communication with the authors, a future version of op. cit. may apply to arbitrary sets of tilings..
@@ 4 Finite local complexity is undecidable for geometric tilings
a shapeset $\mathcal{T}$ of polygonal tiles.
@@
The FLC problem for purely geometric subshifts is $\Sigma_{1}^{0}$-complete and therefore undecidable.
We start by proving the upper bound, namely that it is recursively enumerable.
Lemma 4.3.
The FLC problem is in $\Sigma_{1}^{0}$.
Proof 4.4.
When a geometric subshift does not have FLC, it has tilings with a shear line (see [20]): that is to say that two independent half planes can be tiled independently and shifted continuously relative to each other along this line. In order to prove that FLC is a $\Sigma^{0}_{1}$ property, we will show that having a shear line is a $\Pi_{1}^{0}$ property.
When a shear line appears, then its direction is necessarily the direction of one of the sides of some shape of the input. There is a finite number of such directions. A semi-algorithm to decide the existence of a shear line thus consists in finding if one of these directions cuts the plane in two halves such that each half plane may be perfectly tiled.
As the existence of a tiling for a shapeset is a $\Pi^{0}_{1}$ problem, so is the existence of a tiling of the half plane. The search for tileable half planes along the finite directions of the sides of the shapes may be done concurrently. If no direction can be tiled in perfect half planes, then at some points the finite number of searches will all halt and the algorithm halts, otherwise it runs forever.
Lemma 4.5.
The FLC problem is $\Sigma_{1}^{0}$-hard.
Proof 4.6.
We proceed by reduction to the halting problem for Turing machines. For each Turing machine $M$, we build a shapeset $\mathcal{T}$ such that the purely geometric subshift is not empty and has finite local complexity iff $M$ halts on the empty tape.
We start from $1\times 1$ square tiles decorated with the Robinson [26] decorations as modified by Gurevich and Koryakov [14]: this tileset tiles the plane periodically iff the Turing machine encoded in it halts and aperiodically otherwise, as shown in Figure 12. As we want to effectively embed computations in non-intersecting hierarchical squares of Robinson tilings, we use a 10-tiles variant of the Robinson tileset where odd levels of the hierarchy of squares are blue and even levels of the hierarchy of squares are red. Robinson’s tileset can be defined either by purely geometric or symbolic tilings: see Fig. 13. Gurevich and Koryakov’s tileset can also be made purely geometric: see Figs. 14 and 15.
@@
Figure 13: A purely geometric presentation of the alternating squares variant of Robinson tiles. Decorations are shown for clarity only, and the computing layer is not represented.
@@
(a) Each Robinson or Gurevich-Koryakov computing tile contains 4 computing areas, each area behaving like a Wang tile of the usual Turing machine simulation tileset (see for example [17]).
(b) Schematics of a computing tile where the Robinson or Gurevich-Koryakov layer is abstracted. Each tile simulates two computation steps of a Turing machine.
(c) Two examples of computing tiles. The computing path is represented with thin gray lines and boxes.
Left: starting from state $q_{1}$, the machine reads a 0 and performs two computation steps that result in writing a 0 and going in state $q_{2}$ without moving (the intermediate state and symbol is not specified).
Right: starting from state $q_{2}$, the machine reads a 0, writes a $1$, stays in state $q_{2}$ and moves to the right. Then, the machine again reads a 0, writes a $1$, stays in state $q_{2}$ and moves to the right.
@@
If $M$ does not halt, any individual tiling in the subshift $X$ has finite local complexity (even though $X$ doesn’t), since it contains at most one shear line with a given offset which yields only finitely many $2$-shapes patterns. The construction can be tweaked so that individual tilings do not have FLC in this case by duplicating our shapeset into squares of size 1 and squares of size $\varphi=(1+\sqrt{5})/2$ (or any irrational number). Tiles inside a finite or infinite green square synchronise so that only one size is used. However, when there is a shear line (in the case where the machine does not halt), the two half-planes may use different shapesets, as illustrated on Fig. 19. The offset between adjacent tiles across the shear lines are the multiples of $\varphi$ modulo $1$ (plus an initial offset) which are dense in the interval $[0,1]$. So such a tiling does not have finite local complexity.
@@ Appendix A Aperiodic Wang tilesets from aperiodic geometric tilesets
This construction was first used in [13, §11.1] for Kite-and-Darts Penrose tilings to produce an aperiodic Wang tileset of $32$ tiles, and left as an exercise (11.1.2) for rhombus Penrose tilings to produce an aperiodic Wang tileset of $24$ tiles; full details on this latter case can be found in [16, §5]. We give an overview of this construction, in particular to shed light on geometric and combinatorial conditions that make it work. We give a somewhat formal statement below. We state the result and present the construction for rhombus tilings, though it can be applied to other shapes e.g. the Kites-and-Darts.
Theorem A.1 (folklore).
Let $\mathcal{T}$ be a finite set of rhombus shapes and $\mathbf{T}$ a finite labelled tileset on $\mathcal{T}$. Let $X$ be an edge-to-edge $\mathbf{T}$ nearest-neighbour subshift of finite type, that is, defined by finitely many forbidden patterns made of two edge-adjacent tiles. If there exists a shape $t_{0}\in\mathcal{T}$ such that $t_{0}$ is uniformly recurrent in $X$, then there exists a Wang tileset $\mathbf{T}_{X,t_{0}}$ whose subshift of valid tilings is orbit-equivalent to $X$.
In particular if $X$ is aperiodic, then $\mathbf{T}_{X,t_{0}}$ is an aperiodic Wang tileset.
Moreover if the language of $X$ (at least up to the radius of recurrence of $t_{0}$) can be computed, then $\mathbf{T}_{X,t_{0}}$ can be effectively constructed.
We do not give a precise definition of orbit equivalence [22]. The idea is that there is a well-behaved correspondence between orbits of tilings in $X$ and orbits of valid $\mathbf{T}_{X,t_{0}}$ tilings, which is illustrated in Fig. 23. Note that here we implicitly consider the continuous shift on Wang tilings instead of the usual discrete one.
Let $\mathcal{T}$ be a finite set of rhombus shapes and $\mathbf{T}$ a finite labelled tileset on $\mathcal{T}$. Let $X$ be an edge-to-edge $\mathbf{T}$ nearest-neighbour subshift of finite type, that is, defined by finitely many forbidden patterns made of two edge-adjacent tiles.
@@
  • the size of $t_{0}$-chain-squares is bounded, so there are finitely many different $t_{0}$-chain-squares in $X$. A formal proof of a very similar statement is given in e.g. [15, Lemma 21].
Figure 22: A fragment of labelled Penrose rhombus tiling with occurences of a particular tile highlighted.